### v1

# Recursive regularised lattice Boltzmann method for magnetohydrodynamics

Abstract
We present and test a recursive regularised lattice Boltzmann method for incompressible magnetohydrodynamic (MHD) flows. The approach is based on a double-distribution formulation, in which the magnetic field is evolved using a standard BGK lattice Boltzmann scheme, while the fluid solver is enhanced through a Hermite-based recursive regularisation of the non-equilibrium moments. The method exploits a fourth-order Hermite expansion of the equilibrium distribution on the D2Q9 lattice, allowing higher-order isotropy to be retained while selectively filtering spurious non-hydrodynamic contributions. The regularisation procedure reconstructs the non-equilibrium distribution from physically consistent Hermite coefficients, avoiding explicit evaluation of velocity gradients. The resulting scheme preserves the correct incompressible MHD limit, improves numerical stability at low viscosities, and reduces lattice-dependent artefacts. The proposed formulation provides a robust and versatile framework for MHD simulations and offers a systematic route for extending regularised lattice Boltzmann methods to coupled multiphysics systems.
I Introduction
Magnetohydrodynamics (MHD) governs the coupled dynamics of electrically conducting fluids and magnetic fields and is central to a wide range of physical and technological applications, including plasma physics, astrophysical and geophysical flows, liquid-metal technologies, and electromagnetic flow control biskamp2003mhd. From a modelling standpoint, the incompressible resistive MHD flow is described by the coupled system
$\partial_{t}\mathbf{u}+\mathbf{u}\cdot\nabla\mathbf{u}$ $=-\frac{\nabla p}{\rho_{0}}+\nu\nabla^{2}\mathbf{u}+\frac{(\nabla\times\mathbf{b})\times\mathbf{b}}{\rho_{0}},$ (1)
$\partial_{t}\mathbf{b}+\mathbf{u}\cdot\nabla\mathbf{b}$ $=\mathbf{b}\cdot\nabla\mathbf{u}+\eta\nabla^{2}\mathbf{b},$ (2)
supplemented by the solenoidal constraints
$\nabla\cdot\mathbf{u}=0,\qquad\nabla\cdot\mathbf{b}=0.$ (3)
Here $\partial_{t}$ denotes the partial derivative with respect to time $t$, $\nabla$ is the gradient operator, $\mathbf{u}$ and $\mathbf{b}$ denote the velocity and magnetic fields, respectively, $p$ is the hydrodynamic pressure, $\nu$ is the kinematic viscosity, $\eta$ is magnetic diffusivity, and $\rho_{0}$ is the reference fluid mass density. The numerical prediction of MHD flows is particularly challenging due to the strong non-linear coupling between velocity and magnetic fields, the coexistence of multiple transport mechanisms, and the requirement to satisfy solenoidal constraints on both fields 10.1119/1.1482065. These difficulties are especially pronounced in regimes characterised by low viscosity, low magnetic diffusivity, or strong Lorentz forces, where numerical dissipation and spurious artefacts can significantly affect solution fidelity mininni2006small; FANG2025109400.
The lattice Boltzmann method kruger2016lattice (LBM) has emerged as an attractive mesoscopic alternative to conventional discretisations of the incompressible MHD equations. Briefly, the flow physics arises from the evolution of collections (also known as distributions or populations) of fictitious particles colliding and streaming along the links of a fixed Cartesian lattice. The earliest attempt to construct a lattice gas automaton for magnetohydrodynamics dates back to the work of Montgomery & Doolen MONTGOMERY1987229. Their formulation is based on a magnetic vector potential approach, in which the Lorentz force is incorporated through the evaluation of a Laplacian operator. This requires the introduction of an additional non-local finite-difference procedure. Subsequently, Chen et al. PhysRevLett.67.3776 proposed a purely local lattice gas model. However, this approach requires solving a 36-state MHD cellular automaton at each node of a two-dimensional hexagonal lattice, resulting in a prohibitively high computational cost. Martinez et al. 10.1063/1.870640 later reduced the number of states to 12, partially mitigating this limitation.
An alternative strategy was introduced by Succi et al. PhysRevA.43.4521 who developed a hybrid scheme coupling the lattice Boltzmann method with a finite-difference discretisation for two-dimensional MHD flows. While computationally efficient, this formulation restricts the magnetic Prandtl number,
$\mathrm{Pr_{m}}=\frac{\nu}{\eta},$ (4)
to unity.
A major advance was achieved by Dellar dellar2002lbm_mhd who demonstrated that the MHD equations can be recovered from two coupled lattice Boltzmann equations (LBEs) based on the single-relaxation-time Bhatnagar–Gross–Krook (BGK) collision operator BGK. The first LBE evolves scalar populations $f_{i}$ of fictitious particles carrying mass, from which the fluid density $\rho$ and momentum $\rho\mathbf{u}$ are obtained. The second LBE governs vector-valued populations $\bm{g}_{j}$ that describe the evolution of the magnetic field $\mathbf{b}$. This formulation overcomes the principal limitations of previous approaches: it is fully local, allows for arbitrary magnetic Prandtl numbers, and remains computationally efficient. In the present work, this scheme is adopted as a baseline and extended to develop an improved method tailored to high-Reynolds-number MHD flows.
The numerical properties of lattice Boltzmann MHD models depend strongly on the choice of collision operator. While the BGK model remains widely used due to its simplicity and efficiency, it is well known to suffer from limited numerical stability at low viscosities and from sensitivity to lattice non-hydrodynamic modes latt2006lbgk_reg. In magnetohydrodynamic flows, these limitations are amplified by the Lorentz force and by the transfer of energy across multiple spatial scales DeRosis03062018.
To address these issues, several advanced collision models have been proposed. Multiple-relaxation-time (MRT) formulations based on raw moments (RMs) improve stability by relaxing different kinetic moments at distinct rates lallemand2000theory; dHumieres2002mrt, while lattice Boltzmann methods based on central moments (CMs) and cumulants further enhance robustness by performing the collision process in a locally comoving frame, improving Galilean invariance and isotropy geier2006cascaded; geier2015cm. Central-moment formulations have been successfully applied to a wide range of challenging flow problems, including multiphysics and magnetohydrodynamic systems, demonstrating excellent stability and accuracy DeRosis2017PRE; 10.1063/1.5124719; patnaik2025three.
An alternative stabilisation strategy is provided by regularised lattice Boltzmann methods, in which the non-equilibrium distribution is reconstructed from a truncated Hermite expansion malaspinas2015increasing. More recently, recursive regularisation (RR) techniques have been proposed to systematically reconstruct higher-order non-equilibrium moments from lower-order information, leading to improved isotropy and robustness in hydrodynamic flows coreixas2017recursive; coreixas2020highorder. Despite these advances, including extensions to multiphase flows LU2026105500, recursive regularisation has not yet been systematically explored in the context of magnetohydrodynamic lattice Boltzmann models. Existing work on its application to MHD remains limited in scope and does not provide a comprehensive assessment 10.1007/978-981-95-3401-2_35. The present study fills this gap by offering the first systematic investigation of the method for MHD flows, rigorously assessing accuracy, stability, and scalability over a wide range of regimes. Indeed, we aim to assess whether the stabilisation and isotropy benefits of recursive regularisation observed in hydrodynamic flows carry over to magnetohydrodynamic systems.
To this end, we propose a recursive regularised lattice Boltzmann method for incompressible two-dimensional MHD flows and explicitly contrast its performance with that of three distinct collision models: BGK, MRT and CMs-based LBMs. We adopt a hybrid double-distribution strategy in which the magnetic field is evolved using the BGK scheme proposed by Dellar dellar2002lbm_mhd, while the fluid solver is enhanced through recursive regularisation of the non-equilibrium moments. The approach is based on a fourth-order Hermite expansion of the equilibrium distribution on the D2Q9 lattice and avoids the explicit evaluation of velocity gradients.
The method is validated using the Orszag–Tang vortex, a canonical benchmark for non-linear MHD dynamics featuring strong velocity–magnetic field coupling, current-sheet formation, and a transition toward MHD turbulence orszag1979ot. The impact of regularisation on numerical stability, dissipation, and the accurate capture of key MHD features are assessed.
The paper is organised as follows. Section II recalls the BGK lattice Boltzmann formulation and presents our recursive regularisation procedure. Numerical results for the Orszag–Tang vortex at low and high Reynolds numbers are presented in Section III, followed by concluding remarks in Section IV.
II Methodology
We consider a two-dimensional Cartesian domain with coordinates $(x,y)$, filled with an electrically conducting fluid of reference mass density $\rho_{0}$, kinematic viscosity $\nu$, and magnetic diffusivity $\eta$. Here the gradient operator is $\nabla=[\partial_{x},\partial_{y}]$, and the flow velocity and magnetic field vectors are $\mathbf{u}=[u_{x},u_{y}]$ and $\mathbf{b}=[b_{x},b_{y}]$, respectively. In the remainder of this section, we first briefly recall the BGK lattice Boltzmann formulation for MHD and subsequently introduce our strategy based on recursive regularisation.
II.1 BGK lattice Boltzmann formulation
Following Dellar dellar2002lbm_mhd, the magnetohydrodynamic equations are solved using a double-distribution-function LBM. The fluid dynamics are described by a scalar particle distribution function $f_{i}(\mathbf{x},t)$ defined on a D2Q9 lattice, where the index $i=0,\ldots,8$ labels the discrete velocity directions $\mathbf{c}_{i}=[c_{i,x},c_{i,y}]$, given by latt2007technical
$c_{i,x}$ $=[0,-1,-1,-1,0,1,1,1,0],$ (5)
$c_{i,y}$ $=[0,1,0,-1,-1,-1,0,1,1].$ (6)
$\mathbf{x}$ denotes the spatial location of a given lattice point. The magnetic field is evolved using auxiliary vector-valued distribution functions $\bm{g}_{j}=[g_{j,x},g_{j,y}]$ defined on a D2Q5 lattice, where $j=0,\ldots,4$ and the corresponding lattice velocities $\mathbf{e}_{j}=[e_{j,x},e_{j,y}]$ are
$e_{j,x}=[0,-1,0,1,0],\qquad e_{j,y}=[0,0,-1,0,1].$ (7)
Within the BGK approximation, the LBEs read as follows
$f_{i}(\mathbf{x}+\mathbf{c}_{i}\Delta t,t+\Delta t)=f_{i}^{\star}(\mathbf{x},t),$ (8)
$\bm{g}_{j}(\mathbf{x}+\mathbf{e}_{j}\Delta t,t+\Delta t)=\bm{g}_{j}^{\star}(\mathbf{x},t),$ (9)
where $\Delta t=1$ is the lattice time step and the so-called post-collision populations (denoted by the superscript $\star$) are evaluated as
$f_{i}^{\star}(\mathbf{x},t)=f_{i}^{\mathrm{eq}}(\mathbf{x},t)+\left(1-\omega\right)f_{i}^{\mathrm{neq}}(\mathbf{x},t),$ (10)
$\bm{g}_{j}^{\star}(\mathbf{x},t)=\bm{g}_{j}^{\mathrm{eq}}(\mathbf{x},t)+\left(1-\omega_{\mathrm{m}}\right)\bm{g}_{j}^{\mathrm{neq}}(\mathbf{x},t).$ (11)
The relaxation frequencies $\omega$ and $\omega_{\mathrm{m}}$ are related to the kinematic viscosity $\nu$ and magnetic diffusivity $\eta$ through
$\nu=\left(\frac{1}{\omega}-\frac{1}{2}\right),\qquad\eta=\left(\frac{1}{\omega_{\mathrm{m}}}-\frac{1}{2}\right),$ (12)
respectively. To lighten the notation, the dependence on space and time is implicitly assumed in the following. The equilibrium distribution functions are given by
$f_{i}^{\mathrm{eq}}$ $=w_{i}\rho\left[1+\frac{\mathcal{H}_{i}^{(1)}\cdot\mathbf{u}}{c_{s}^{2}}+\frac{1}{2c_{s}^{4}}\mathcal{H}_{i}^{(2)}:\mathbf{u}\mathbf{u}\right]+$
$\quad\frac{w_{i}}{2c_{s}^{4}}\left[\frac{1}{2}|\mathbf{b}|^{2}|\mathbf{c}_{i}|^{2}-(\mathbf{c}_{i}\cdot\mathbf{b})^{2}\right],$ (13)
$g_{j,\alpha}^{\mathrm{eq}}$ $=\psi_{j}\left[b_{\alpha}+\frac{e_{j,\beta}}{c_{s}^{2}}\left(u_{\beta}b_{\alpha}-u_{\alpha}b_{\beta}\right)\right],$ (14)
where $c_{s}=1/\sqrt{3}$ is the lattice sound speed. The weighting factors for the velocity field are
$w_{0}=\frac{4}{9},\qquad w_{1,3,5,7}=\frac{1}{36},\qquad w_{2,4,6,8}=\frac{1}{9},$ (15)
while for the magnetic field
$\psi_{0}=\frac{1}{3},\qquad\psi_{1,2,3,4}=\frac{1}{6}.$ (16)
The non-equilibrium terms are simply computed as
$f_{i}^{\mathrm{neq}}$ $=$ $f_{i}-f_{i}^{\mathrm{eq}},$ (17)
$\bm{g}_{j}^{\mathrm{neq}}$ $=$ $\bm{g}_{j}-\bm{g}_{j}^{\mathrm{eq}}.$ (18)
Rather than using the second-order equilibrium (13), we follow the approach of Malaspinas malaspinas2015increasing and Coreixas et al. coreixas2017recursive, who showed that the full isotropy of the D2Q9 lattice can be exploited by expanding the equilibrium distribution in a Hermite basis up to fourth order. The resulting equilibrium reads
$f_{i}^{\mathrm{eq}}=w_{i}\rho\Bigg[$ $\underbrace{1}_{\mathcal{O}(\mathbf{u}^{0})}+\underbrace{\frac{\mathcal{H}_{i}^{(1)}\cdot\mathbf{u}}{c_{s}^{2}}}_{\mathcal{O}(\mathbf{u}^{1})}+\underbrace{\frac{1}{2c_{s}^{4}}\mathcal{H}_{i}^{(2)}:\mathbf{u}\mathbf{u}}_{\mathcal{O}(\mathbf{u}^{2})}+$
$\underbrace{\frac{1}{2c_{s}^{6}}\left(\mathcal{H}^{(3)}_{i,xxy}u_{x}^{2}u_{y}+\mathcal{H}^{(3)}_{i,xyy}u_{x}u_{y}^{2}\right)}_{\mathcal{O}(\mathbf{u}^{3})}+$
$\underbrace{\mathcal{H}^{(4)}_{i,xxyy}u_{x}^{2}u_{y}^{2}}_{\mathcal{O}(\mathbf{u}^{4})}\Bigg]+$
$\frac{w_{i}}{2c_{s}^{4}}\underbrace{\left[\frac{1}{2}|\mathbf{b}|^{2}|\mathbf{c}_{i}|^{2}-(\mathbf{c}_{i}\cdot\mathbf{b})^{2}\right]}_{\text{Maxwell stress }\mathcal{O}(\mathbf{b}^{2})}.$ (19)
The Hermite tensors are defined as
$\mathcal{H}_{i}^{(0)}$ $=1,$
$\mathcal{H}_{i,\alpha}^{(1)}$ $=c_{i,\alpha},$
$\mathcal{H}_{i,\alpha\beta}^{(2)}$ $=c_{i,\alpha}c_{i,\beta}-c_{s}^{2}\delta_{\alpha\beta},$
$\mathcal{H}_{i,\alpha\beta\gamma}^{(3)}$ $=c_{i,\alpha}c_{i,\beta}c_{i,\gamma}-c_{s}^{2}\left(c_{i,\alpha}\delta_{\beta\gamma}+c_{i,\beta}\delta_{\alpha\gamma}+c_{i,\gamma}\delta_{\alpha\beta}\right),$
$\mathcal{H}_{i,\alpha\beta\gamma\delta}^{(4)}$ $=c_{i,\alpha}c_{i,\beta}c_{i,\gamma}c_{i,\delta}-c_{s}^{2}\sum_{\text{perm}}c_{i,\alpha}c_{i,\beta}\delta_{\gamma\delta}+$
$c_{s}^{4}\sum_{\text{perm}}\delta_{\alpha\beta}\delta_{\gamma\delta},$ (20)
where $\delta_{\alpha\beta}$ is the Kronecker delta and $(\alpha,\beta,\gamma,\delta)\in\{x,y\}^{4}$. The equilibrium distribution $f_{i}^{\mathrm{eq}}$ is fixed throughout this study to the fourth-order Hermite formulation given in Eq. (19).
Finally, the macroscopic fields are obtained as
$\rho=\sum_{i}f_{i},\qquad\mathbf{u}=\frac{\sum_{i}f_{i}\mathbf{c}_{i}}{\rho},\qquad\mathbf{b}=\sum_{j}\bm{g}_{j}.$ (21)
It is worth noting that the LBM allows us to compute the electric current $j_{z}=\nabla\times\mathbf{b}$ locally from the populations as PATTISON2008557
$j_{z}=-\frac{\omega_{\mathrm{m}}}{c_{s}^{2}}\varepsilon_{\alpha\beta}\sum_{j}\left(e_{j,\alpha}g_{j,\beta}-e_{j,\alpha}g_{j,\beta}^{\mathrm{eq}}\right),$ (22)
where
$\sum_{j}e_{j,\alpha}g_{j,\beta}^{\mathrm{eq}}=u_{\beta}b_{\alpha}-u_{\alpha}b_{\beta},$ (23)
and $\varepsilon_{\alpha\beta}$ is the Levi-Civita permutation tensor, without the need to evaluate non-local finite-difference operators.
II.2 Recursive regularised lattice Boltzmann method
We now derive a recursive regularisation strategy for magnetohydrodynamic flows. Specifically, we adopt a hybrid approach in which the magnetic field is evolved using the standard BGK lattice Boltzmann scheme described above, while a dedicated regularisation procedure is applied to the fluid solver in order to enhance numerical stability and accuracy.
Following Coreixas et al. coreixas2017recursive, we can express the post-collision state $f_{i}^{\star}$ in terms of a truncated Hermite expansion as
[…]
The equilibrium Hermite coefficients $\bm{a}_{0}^{(n)}$ are given by
$a_{0}^{(0)}$ $=\rho,$
$a_{0,\alpha}^{(1)}$ $=\rho u_{\alpha},$
$a_{0,\alpha\beta}^{(2)}$ $=\rho u_{\alpha}u_{\beta}+M_{\alpha\beta},$
$a^{(3)}_{0,\alpha\beta\gamma}$ $=\rho\,u_{\alpha}u_{\beta}u_{\gamma}+u_{\alpha}M_{\beta\gamma}+u_{\beta}M_{\alpha\gamma}+u_{\gamma}M_{\alpha\beta},$
$a^{(4)}_{0,\alpha\beta\gamma\delta}$ $=\rho\,u_{\alpha}u_{\beta}u_{\gamma}u_{\delta}+\Big(u_{\alpha}u_{\beta}M_{\gamma\delta}+u_{\alpha}u_{\gamma}M_{\beta\delta}+$
$\quad u_{\alpha}u_{\delta}M_{\beta\gamma}+u_{\beta}u_{\gamma}M_{\alpha\delta}+$
$\quad u_{\beta}u_{\delta}M_{\alpha\gamma}+u_{\gamma}u_{\delta}M_{\alpha\beta}\Big),$ (25)
where the Maxwell stress tensor is defined as dellar2002lbm_mhd
$M_{\alpha\beta}=\frac{1}{2}|\mathbf{b}|^{2}\delta_{\alpha\beta}-b_{\alpha}b_{\beta}.$ (26)
When the magnetic field is neglected, Eqs. (25) reduce exactly to the pure hydrodynamic formulation coreixas2017recursive.
The corresponding non-equilibrium Hermite coefficients $\bm{a}_{1}^{(n)}$ read
$a_{1}^{(0)}$ $=0,$
$a_{1,\alpha}^{(1)}$ $=0,$
$a_{1,\alpha\beta}^{(2)}$ $=-\nu\rho c_{s}^{2}\left[\partial_{\alpha}u_{\beta}+\partial_{\beta}u_{\alpha}-(\partial_{\gamma}u_{\gamma})\delta_{\alpha\beta}\right],$
$a_{1,\alpha\beta\gamma}^{(3)}$ $=u_{\alpha}a_{1,\beta\gamma}^{(2)}+u_{\beta}a_{1,\alpha\gamma}^{(2)}+u_{\gamma}a_{1,\alpha\beta}^{(2)},$
$a_{1,\alpha\beta\gamma\delta}^{(4)}$ $=u_{\alpha}a_{1,\beta\gamma\delta}^{(3)}+u_{\beta}a_{1,\alpha\gamma\delta}^{(3)}+u_{\gamma}a_{1,\alpha\beta\delta}^{(3)}+u_{\delta}a_{1,\alpha\beta\gamma}^{(3)}$-
$\Big[u_{\alpha}u_{\beta}a_{1,\gamma\delta}^{(2)}+u_{\alpha}u_{\gamma}a_{1,\beta\delta}^{(2)}+u_{\alpha}u_{\delta}a_{1,\beta\gamma}^{(2)}+$
$u_{\beta}u_{\gamma}a_{1,\alpha\delta}^{(2)}+u_{\beta}u_{\delta}a_{1,\alpha\gamma}^{(2)}+u_{\gamma}u_{\delta}a_{1,\alpha\beta}^{(2)}\Big].$ (27)
Rather than explicitly computing spatial derivatives of the velocity field, the second-order non-equilibrium Hermite coefficient $a_{1,\alpha\beta}^{(2)}$ is approximated directly from the distribution functions as
$a_{1,\alpha\beta}^{(2)}\approx\sum_{i}\mathcal{H}_{i,\alpha\beta}^{(2)}\left(f_{i}-f_{i}^{\mathrm{eq}}\right).$ (28)
It is worth noting that Eq. (24) is formally equivalent to
$f_{i}^{\star}=f_{i}^{\mathrm{eq}}+(1-\omega)f_{i}^{\mathrm{neq,reg}},$ (29)
with
[…]
Here, the former equilibrium contribution is given by Eq. (19), while the latter regularised non-equilibrium counterpart is
[…]
which becomes
$f_{i}^{\mathrm{neq,reg}}=w_{i}\left[\frac{1}{2c_{s}^{4}}a^{(2)}_{1,\alpha\beta}\mathcal{H}^{(2)}_{i,\alpha\beta}+\frac{1}{6c_{s}^{6}}a^{(3)}_{1,\alpha\beta\gamma}\mathcal{H}^{(3)}_{i,\alpha\beta\gamma}+\frac{1}{24c_{s}^{8}}a^{(4)}_{1,\alpha\beta\gamma\delta}\mathcal{H}^{(4)}_{i,\alpha\beta\gamma\delta}\right$]. (33)
Algorithm of computation
Within a single time step, the proposed procedure executes the following six tasks at each grid point.
  1. 1.
    Predict the macroscopic variables:
    $\rho=\sum_{i}f_{i},\qquad\mathbf{u}=\frac{\sum_{i}f_{i}\mathbf{c}_{i}}{\rho},\qquad\mathbf{b}=\sum_{j}\bm{g}_{j}.$
  2. 2.
    Evaluate the equilibrium populations:
    $f_{i}^{\mathrm{eq}}=w_{i}\rho\Bigg[$ $1+\frac{\mathcal{H}_{i}^{(1)}\cdot\mathbf{u}}{c_{s}^{2}}+\frac{1}{2c_{s}^{4}}\mathcal{H}_{i}^{(2)}:\mathbf{u}\mathbf{u}+$
    $\frac{1}{2c_{s}^{6}}\left(\mathcal{H}^{(3)}_{i,xxy}u_{x}^{2}u_{y}+\mathcal{H}^{(3)}_{i,xyy}u_{x}u_{y}^{2}\right)+$
    $\mathcal{H}^{(4)}_{i,xxyy}u_{x}^{2}u_{y}^{2}\Bigg]+$
    $\frac{w_{i}}{2c_{s}^{4}}\left[\frac{1}{2}|\mathbf{b}|^{2}|\mathbf{c}_{i}|^{2}-(\mathbf{c}_{i}\cdot\mathbf{b})^{2}\right],$
    $g_{j,\alpha}^{\mathrm{eq}}$ $=\psi_{j}\left[b_{\alpha}+\frac{e_{j,\beta}}{c_{s}^{2}}\left(u_{\beta}b_{\alpha}-u_{\alpha}b_{\beta}\right)\right].$
  3. 3.
    Compute the non-equilibrium Hermite coefficients:
    $a_{1,\alpha\beta}^{(2)}$ $=\sum_{i}\mathcal{H}_{i,\alpha\beta}^{(2)}\left(f_{i}-f_{i}^{\mathrm{eq}}\right),$
    $a_{1,\alpha\beta\gamma}^{(3)}$ $=u_{\alpha}a_{1,\beta\gamma}^{(2)}+u_{\beta}a_{1,\alpha\gamma}^{(2)}+u_{\gamma}a_{1,\alpha\beta}^{(2)},$
    $a_{1,\alpha\beta\gamma\delta}^{(4)}$ $=u_{\alpha}a_{1,\beta\gamma\delta}^{(3)}+u_{\beta}a_{1,\alpha\gamma\delta}^{(3)}+u_{\gamma}a_{1,\alpha\beta\delta}^{(3)}+u_{\delta}a_{1,\alpha\beta\gamma}^{(3)}$-
    $\quad\Big[u_{\alpha}u_{\beta}a_{1,\gamma\delta}^{(2)}+u_{\alpha}u_{\gamma}a_{1,\beta\delta}^{(2)}+u_{\alpha}u_{\delta}a_{1,\beta\gamma}^{(2)}+$
    $\quad u_{\beta}u_{\gamma}a_{1,\alpha\delta}^{(2)}+u_{\beta}u_{\delta}a_{1,\alpha\gamma}^{(2)}+u_{\gamma}u_{\delta}a_{1,\alpha\beta}^{(2)}\Big].$
  4. 4.
    Reconstruct the non-equilibrium regularised populations:
[…]
    $\left.\frac{1}{24c_{s}^{8}}a^{(4)}_{1,\alpha\beta\gamma\delta}\mathcal{H}^{(4)}_{i,\alpha\beta\gamma\delta}\right$].
  5. 5.
    Perform the collision step:
    $f_{i}^{\star}=f_{i}^{\mathrm{eq}}+\left(1-\omega\right)f_{i}^{\mathrm{neq,req}},$
    $\bm{g}_{j}^{\star}=\bm{g}_{j}^{\mathrm{eq}}+\left(1-\omega_{\mathrm{m}}\right)\left(\bm{g}_{j}-\bm{g}_{j}^{\mathrm{eq}}\right).$
  6. 6.
    Stream by
    $f_{i}(\mathbf{x}+\mathbf{c}_{i}\Delta t,t+\Delta t)$ $=f_{i}^{\star}(\mathbf{x},t),$
    $\bm{g}_{j}(\mathbf{x}+\mathbf{e}_{j}\Delta t,t+\Delta t)$ $=\bm{g}_{j}^{\star}(\mathbf{x},t),$
    and advance in time.
It is worth noting that the streaming step is implemented using the swap technique proposed by Latt latt2007technical, which enables the storage of only a single copy of the particle distribution functions. The interested reader is referred to the works of Geier & Schönherr computation5020019 and Lehmann computation10060092 for more recent approaches that likewise achieve single-population storage while preserving computational efficiency.
III Orszag–Tang vortex
We assess the proposed recursive regularised formulation using the two-dimensional Orszag–Tang vortex, a canonical benchmark for non-linear MHD dynamics that exhibits rapid current-sheet formation and a subsequent transition toward MHD turbulence orszag1979ot.
The initial velocity and magnetic fields are prescribed as
$\mathbf{u}(\mathbf{x},t=0)$ $=$ $u_{0}\left[-\sin(y),\,\sin(x)\right],$
$\mathbf{b}(\mathbf{x},t=0)$ $=$ $b_{0}\left[-\sin(y),\,\sin(2x)\right],$ (34)
which satisfy the solenoidal constraints $\nabla\cdot\mathbf{u}=0$ and $\nabla\cdot\mathbf{b}=0$ by construction. The fluid mass density is initialised uniformly as $\rho_{0}=1$. All simulations are performed in a doubly periodic square domain $\Omega=[0,2\pi]\times[0,2\pi]$, discretised using a uniform Cartesian grid of $N\times N$ lattice nodes. The flow is characterised by the Reynolds and magnetic Prandtl numbers, with the former being defined as
$\mathrm{Re}=\frac{u_{0}L}{\nu},$ (35)
where $L=2\pi$ denotes the characteristic length scale, $\nu$ is the kinematic viscosity, and $u_{0}=b_{0}=2$ being the root-mean-square of the fields in Eqs. (34). The Mach number of our simulations is set to $\mathrm{Ma}=\tilde{u}_{0}/c_{s}\simeq 0.009$, where $\tilde{u}_{0}=0.2\,\pi^{-1}$ is the reference velocity expressed in lattice units. This value is sufficiently small to suppress spurious compressibility effects in the solution of the lattice Boltzmann equations. The choice of $\tilde{u}_{0}$ as a multiple of $\pi^{-1}$ further facilitates sampling of flow quantities at salient discrete time steps.
Spatial derivatives are consistently evaluated using the isotropic lattice gradient
$\partial_{\alpha}u_{\beta}(\mathbf{x})$ $=$ $c_{s}^{-2}\sum_{i}\left[w_{i}u_{\beta}(\mathbf{x}+\mathbf{c}_{i})c_{i,\alpha}\right],$
$\partial_{\alpha}b_{\beta}(\mathbf{x})$ $=$ $c_{s}^{-2}\sum_{i}\left[w_{i}b_{\beta}(\mathbf{x}+\mathbf{c}_{i})c_{i,\alpha}\right],$ (36)
which is fully consistent with the underlying lattice symmetry https://doi.org/10.1155/2014/142907; PhysRevE.96.013317; PhysRevE.98.013305.
We first consider a low-Reynolds-number configuration and benchmark our results against the spectral reference solution of Ref. dellar2002lbm_mhd, as well as against three alternative lattice Boltzmann collision operators, the BGK, raw-moments-based, and central-moments-based formulations, which all employ fourth-order truncated equilibria for $f_{i}$ in order to ensure a fair assessment. We then extend the analysis to turbulent regimes. Eventually, the computational cost involved by the present procedure is discussed.
III.1 Low-Reynolds-number flow
Following Orszag & Tang orszag1979ot and Dellar dellar2002lbm_mhd, we consider the case $\mathrm{Re}=200\pi$ and monitor the temporal evolution of the peak electric current density, $j_{\max}=\max_{\Omega}|j_{z}|$, and peak vorticity, $\omega_{\max}=\max_{\Omega}|\omega_{z}|$, where $j_{z}$ is evaluated from Eq. (22) and $\omega_{z}=\nabla\times\mathbf{u}=\partial_{x}u_{y}-\partial_{y}u_{x}$ by finite differences. All simulations employ identical physical parameters and initial conditions in order to isolate the effect of the collision operator. The magnetic Prandtl number is fixed to unity.
Tables 1 and 2 report the values of $j_{\max}$ and $\omega_{\max}$ at representative times across progressively refined grids, together with the spectral reference solution.
At early times ($t=0.5$), the flow remains smooth and well resolved. In this regime, all lattice Boltzmann formulations exhibit excellent agreement with the reference data. For the current density, the relative error is approximately $1.5\%$ at $N=128$ and decreases monotonically to $0.66\%$ at $N=512$ for all schemes. Similarly, for the vorticity, the finest-grid error drops below $5\times 10^{-4}$, indicating that when gradients are mild the choice of collision model has negligible influence on accuracy.
As the system evolves toward $t=1.0$, sharp magnetic gradients and emerging current sheets substantially increase the numerical stiffness of the problem. On the coarsest grid ($N=128$), all LBMs underpredict both extrema. The BGK scheme yields slightly smaller discrepancies, with relative differences of $3.94\%$ for $j_{\max}$ and $10.09\%$ for $\omega_{\max}$, while the recursive regularised model displays marginally higher dissipation, reaching $4.23\%$ and $10.95\%$, respectively. This behaviour indicates that regularisation introduces a modest additional damping of under-resolved non-equilibrium contributions.
Importantly, these differences vanish rapidly under grid refinement. At $N=512$, all collision models converge toward the spectral reference with relative discrepancies of approximately $1.3\%$ for $j_{\max}$ and $1.4\%$ for $\omega_{\max}$. This confirms that while recursive regularisation is slightly more dissipative on coarse meshes, it remains fully consistent with the reference solution and accurately captures the late-time nonlinear dynamics once sufficient resolution is provided.
$t$ NN Ref. dellar2002lbm_mhd BGK RMs CMs $R$
0.5 128 17.96 17.69 17.69 17.69 17.69
256 18.20 17.98 17.98 17.98 17.98
512 18.24 18.12 18.12 18.12 18.12
1.0 128 45.13 43.35 43.25 43.24 43.22
256 46.30 45.21 45.19 45.19 45.19
512 46.59 45.98 45.98 45.97 45.97
Table 1: Low-Reynolds-number flow: maximum current density $j_{\max}$ at times $t=0.5$ and $t=1.0$, computed on progressively refined grids. Results obtained with different collision models are compared against a spectral reference solution.
$t$ NN Ref. dellar2002lbm_mhd BGK RMs CMs $R$
0.5 128 6.744 6.670 6.666 6.666 6.670
256 6.754 6.737 6.734 6.735 6.737
512 6.758 6.755 6.753 6.753 6.755
1.0 128 14.07 12.65 12.57 12.56 12.53
256 14.16 13.65 13.64 13.64 13.61
512 14.20 14.01 14.00 14.00 13.99
Table 2: Low-Reynolds-number flow: maximum vorticity $\omega_{\max}$ at times $t=0.5$ and $t=1.0$, computed on progressively refined grids. Results obtained with different collision models are compared against a spectral reference solution.
The preservation of the solenoidal constraint, $\nabla\cdot\mathbf{b}=0$, is essential for physical fidelity in MHD simulations. Table 3 reports the maximum magnetic-field divergence. At $t=1.0$, the divergence decreases from approximately $4.6\times 10^{-1}$ at $N=128$ to $4.0\times 10^{-2}$ at $N=512$ for all lattice Boltzmann schemes. Notably, the LBM results remain slightly below the spectral reference values. The near-identical behaviour across collision models demonstrates that divergence control is primarily governed by the discrete magnetic-field representation rather than by the specific regularisation of non-equilibrium moments.
$t$ NN Ref. dellar2002lbm_mhd BGK RMs CMs $R$
0.5 128 0.0939 0.0619 0.0620 0.0620 0.0620
256 0.0246 0.0201 0.0201 0.0201 0.0201
512 0.0062 0.0056 0.0056 0.0056 0.0056
1.0 128 0.5642 0.4623 0.4583 0.4584 0.4587
256 0.1611 0.1513 0.1512 0.1512 0.1512
512 0.0415 0.0402 0.0402 0.0402 0.0402
Table 3: Low-Reynolds-number flow: maximum magnetic-field divergence $\max_{\Omega}|\nabla\cdot\mathbf{b}|$ at times $t=0.5$ and $t=1.0$, computed on progressively refined grids. Results obtained with different collision models are compared against a spectral reference solution.
The sequence shown in Fig. 1 illustrates the temporal evolution of the magnetic field obtained by our present RR procedure. Initially, the field retains the smooth, large-scale structure imposed by the initial condition, with well-organised alternating regions of magnetic intensity. As non-linear coupling intensifies, magnetic field lines undergo progressive stretching and folding, leading to the formation of thin filaments and elongated gradients. These structures correspond to the emergence of current sheets aligned with evolving shear layers, signalling the onset of magnetic reconnection. At later times, the magnetic topology becomes increasingly intermittent and complex, accompanied by a breakdown of the initial symmetries. This evolution directly underlies the growth of $j_{\max}$ and $\omega_{\max}$ observed in the tables and reflects the transition toward fully non-linear MHD dynamics. Moreover, it corroborates the observations previously obtained by other collision operators dellar2002lbm_mhd; DeRosis03062018; DeRosis2017PRE; 10.1063/5.0058884.
(a) $t=0.2$
(b) $t=0.3$
(c) $t=0.4$
(d) $t=0.5$
(e) $t=0.6$
(f) $t=0.7$
(g) $t=0.8$
(h) $t=0.9$
(i) $t=1.0$
Figure 1: Low-Reynolds-number flow: evolution of the magnetic field at selected time instants obtained by the RR LBM. The sequence illustrates the progressive distortion of the initial large-scale structures, the formation of thin current sheets, and the emergence of a highly intermittent magnetic field as the flow transitions toward fully non-linear MHD dynamics.
III.2 Turbulent regimes
We next investigate strongly non-linear regimes corresponding to $\mathrm{Re}=2500$ and $5000$. Tables 47 report the evolution of the peak current density and vorticity across progressively refined grids. These quantities directly probe the formation of thin current sheets and intense shear layers associated with magnetic reconnection and the ensuing turbulent cascade.
At early time ($t=0.5$), the flow remains smooth and weakly non-linear. All collision models yield indistinguishable results and exhibit rapid grid convergence. For instance, at $\mathrm{Re}=2500$, variations in $j_{\max}$ between $N=500$ and $N=2000$ remain below $1\%$, indicating that modest resolutions are sufficient prior to current-sheet formation.
As reconnection develops ($t\gtrsim 1$), sharp gradients emerge and the standard BGK formulation becomes unstable on coarse grids, systematically failing as current concentrations intensify. This breakdown coincides with the onset of thin current sheets and strong vortex roll-up, highlighting the inability of the BGK operator to control high-order non-equilibrium moments in severely under-resolved regimes. In contrast, all regularised and moment-based schemes remain fully stable throughout the evolution. Between $N=1000$ and $N=2000$, relative variations remain below $3\times 10^{-3}$ for $j_{\max}$ and below $5\times 10^{-3}$ for $\omega_{\max}$ at all times, demonstrating very good grid convergence even in strongly turbulent conditions.
Interestingly, the monotonic growth of $j_{\max}$ up to $t\approx 1$ reflects the progressive thinning of current sheets as magnetic field lines collapse and intensify. The subsequent decrease corresponds to reconnection events that release magnetic energy and promote turbulent mixing. Simultaneously, the growth of $\omega_{\max}$ indicates strong vortex stretching around reconnecting structures, confirming the tight coupling between magnetic reconnection and shear-layer formation that drives the turbulent cascade.
Across both turbulent regimes, the recursive regularisation consistently reproduces the converged extrema with accuracy comparable to moment-based formulations while preserving stability even on the coarsest grids. This demonstrates that recursive filtering of non-equilibrium moments is sufficient to control under-resolved dynamics and accurately capture reconnection-driven intermittency, without requiring explicit moment reconstruction.
$t$ NN BGK RMs CMs $R$
0.5 500 22.68 22.67 22.67 22.68
1000 22.82 22.83 22.83 22.83
2000 22.90 22.90 22.90 22.90
1.0 500 Fail 111.46 111.46 111.48
1000 111.40 111.40 111.40 111.41
2000 111.37 111.38 111.38 111.38
1.5 500 Fail 94.75 94.75 94.75
1000 94.43 94.44 94.44 94.44
2000 94.36 94.37 94.37 94.37
2.0 500 Fail 82.61 82.61 82.55
1000 82.41 82.45 82.45 82.42
2000 82.37 82.38 82.38 82.37
2.5 500 Fail 49.97 49.97 49.96
1000 50.33 50.33 50.33 50.33
2000 50.41 50.41 50.41 50.41
3.0 500 Fail 66.20 66.20 66.19
1000 66.08 66.09 66.09 66.08
2000 66.02 66.03 66.03 66.03
3.5 500 Fail 61.57 61.57 61.53
1000 62.26 62.27 62.27 62.26
2000 62.11 62.12 62.12 62.11
4.0 500 Fail 58.56 58.55 58.61
1000 58.64 58.64 58.64 58.65
2000 58.56 58.58 58.58 58.57
Table 4: $\mathrm{Re}=2500$: maximum current density $j_{\max}$ at representative time instants, computed on progressively refined grids.
$t$ NN BGK RMs CMs $R$
0.5 500 7.408 7.408 7.408 7.408
1000 7.561 7.561 7.561 7.561
2000 7.652 7.652 7.652 7.652
1.0 500 Fail 33.58 33.44 33.44
1000 34.79 34.78 34.80 34.78
2000 35.18 35.18 35.19 35.18
1.5 500 Fail 31.06 31.01 31.02
1000 31.93 31.92 31.91 31.92
2000 32.14 32.13 32.13 32.13
2.0 500 Fail 36.42 36.28 36.32
1000 37.44 37.44 37.43 37.46
2000 37.69 37.68 37.68 37.69
2.5 500 Fail 26.13 26.13 26.15
1000 26.97 26.96 26.96 26.97
2000 27.21 27.21 27.21 27.21
3.0 500 Fail 27.70 27.70 27.66
1000 28.44 28.45 28.45 28.43
2000 28.65 28.66 28.66 28.65
3.5 500 Fail 27.66 27.65 27.64
1000 28.27 28.28 28.28 28.28
2000 28.38 28.39 28.38 28.38
4.0 500 Fail 25.30 25.29 25.25
1000 25.92 25.93 25.93 25.91
2000 26.10 26.11 26.11 26.10
Table 5: $\mathrm{Re}=2500$: maximum vorticity $\omega_{\max}$ at representative time instants, computed on progressively refined grids.
$t$ NN BGK RMs CMs $R$
0.5 500 Fail 23.71 23.71 23.72
1000 469.97 23.90 23.90 23.90
2000 23.98 23.98 23.98 23.98
1.0 500 Fail 168.97 169.00 169.06
1000 Fail 169.66 169.66 169.69
2000 169.67 169.67 169.67 169.69
1.5 500 Fail 144.32 144.31 144.32
1000 Fail 142.95 142.95 142.96
2000 142.53 142.54 142.54 142.55
2.0 500 Fail 131.43 131.4 131.33
1000 Fail 131.53 131.52 131.49
2000 131.42 131.45 131.45 131.43
2.5 500 Fail 93.58 93.58 93.60
1000 Fail 93.41 93.41 93.45
2000 92.94 92.92 92.92 92.94
3.0 500 Fail 79.32 79.26 79.96
1000 Fail 76.10 76.09 76.50
2000 75.58 75.58 75.58 75.59
3.5 500 Fail 201.32 201.62 202.14
1000 Fail 190.36 190.36 193.03
2000 180.12 180.15 180.15 182.47
4.0 500 Fail 129.37 129.33 129.22
1000 Fail 130.67 130.66 130.53
2000 130.99 131.04 131.03 130.94
Table 6: $\mathrm{Re}=5000$: maximum current density $j_{\max}$ at representative time instants, computed on progressively refined grids.
$t$ NN BGK RMs CMs $R$
0.5 500 Fail 7.698 7.698 7.697
1000 46.2653 7.808 7.808 7.807
2000 7.886 7.886 7.886 7.886
1.0 500 Fail 49.65 48.96 48.98
1000 Fail 51.75 51.64 51.60
2000 52.61 52.61 52.62 52.60
1.5 500 Fail 47.49 47.49 47.28
1000 Fail 49.31 49.27 49.27
2000 49.82 49.80 49.79 49.80
2.0 500 Fail 57.69 56.85 57.01
1000 Fail 60.54 60.41 60.45
2000 61.13 61.13 61.12 61.14
2.5 500 Fail 45.88 45.82 45.52
1000 Fail 47.57 47.56 47.39
2000 47.99 48.01 48.01 47.92
3.0 500 Fail 45.35 45.41 45.38
1000 Fail 47.16 47.17 47.11
2000 47.51 47.51 47.52 47.47
3.5 500 Fail 66.72 66.45 65.90
1000 Fail 74.40 74.40 74.90
2000 75.34 75.33 75.34 75.84
4.0 500 Fail 49.42 49.32 49.82
1000 Fail 53.50 53.48 53.76
2000 54.51 54.53 54.53 54.67
Table 7: $\mathrm{Re}=5000$: maximum vorticity $\omega_{\max}$ at representative time instants, computed on progressively refined grids.
Fig. 2 depicts the temporal evolution of the magnetic field magnitude in the Orszag–Tang vortex at $\mathrm{Re}=500$, revealing the progressive development of MHD turbulence through field line stretching, sheet formation, and topological rearrangement.
At early times (t=0.5t=0.51.01.0), the magnetic structures remain smooth and large-scale, closely following the organised vortex pattern imposed by the initial condition. As non-linear coupling between the velocity and magnetic fields intensifies (t=1.5t=1.52.02.0), the magnetic field undergoes strong elongation and folding, leading to the formation of narrow regions of enhanced gradients. These localised structures mark the onset of current sheet formation, where magnetic energy is concentrated and dissipative processes become increasingly relevant.
In the later stages (t=2.5t=2.54.04.0), the flow transitions toward a fully developed turbulent regime characterized by thin, intermittent magnetic filaments and sheet-like structures distributed throughout the domain. Such features are classical signatures of magnetic reconnection events, whereby distorted field lines rapidly change topology and release stored magnetic energy across multiple scales. The coexistence of intense localised structures with broader turbulent patterns reflects the underlying cascade of magnetic energy from large to small scales.
The accurate capture of sharp current sheets and reconnection-driven dynamics without spurious oscillations or excessive diffusion highlights the robustness of the present formulation in resolving multiscale MHD phenomena at high Reynolds numbers.
(a) $t=0.5$
(b) $t=1.0$
(c) $t=1.5$
(d) $t=2.0$
(e) $t=2.5$
(f) $t=3.0$
(g) $t=3.5$
(h) $t=4.0$
Figure 2: $\mathrm{Re}=5000$: evolution of the magnetic field at selected time instants. Snapshots illustrate the transition from large-scale coherent structures to fully developed turbulent magnetic filaments.
Synthesis across flow regimes
Taken together, the low-Reynolds-number and turbulent results provide a coherent picture of the role of non-equilibrium control in MHD lattice Boltzmann models. In mildly non-linear regimes dominated by smooth structures, all collision operators perform equivalently, confirming that regularisation does not degrade accuracy in resolved flows. As the dynamics transition toward strongly non-linear behaviour characterised by thin current sheets, magnetic reconnection, and intense shear layers, the numerical challenge shifts from accuracy to stability and robustness. In this regime, the failure of the BGK scheme contrasts sharply with the sustained convergence of more sophisticated formulations. Recursive regularisation emerges as a strategy that preserves accuracy in laminar conditions while ensuring stability in turbulent MHD flows, enabling reliable simulation across the full spectrum of reconnection-driven dynamics encountered in the Orszag–Tang vortex.
Fig. 3 reports the temporal evolution of the maximum electric current density. This quantity directly probes the formation and intensification of current sheets and therefore provides a sensitive measure of magnetic reconnection and non-linear MHD activity. At early times ($t\lesssim 0.5$), all curves collapse closely, indicating that the flow remains smooth and weakly non-linear, with magnetic gradients growing in a similar manner regardless of Reynolds number. During this phase, current amplification is dominated by large-scale field-line stretching imposed by the initial vortex configuration. As the system evolves toward $t\approx 1$, a rapid increase in $j_{\max}$ is observed, corresponding to the progressive thinning of current sheets and the onset of reconnection. This growth becomes increasingly pronounced with Reynolds number. While the low-Re case exhibits a relatively mild peak, both turbulent regimes display substantially stronger current intensification, reflecting reduced dissipation and sharper magnetic gradients. The ordering $j_{\max}(\mathrm{Re}=5000)>j_{\max}(\mathrm{Re}=2500)\gg j_{\max}(\mathrm{Re}=200\pi)$ highlights the enhanced collapse of current layers as inertial effects dominate over resistive smoothing. Beyond the peak, $j_{\max}$ decreases as reconnection events release magnetic energy and promote topological reorganisation of the field. This decay is more gradual and occurs at higher levels in the turbulent regimes, consistent with sustained intermittency and the continuous formation of secondary current sheets characteristic of fully developed MHD turbulence. In contrast, the low-Re flow rapidly relaxes toward a smoother state with significantly weaker magnetic gradients.
Figure 3: Temporal evolution of the peak electric current density $j_{\max}=\max_{\Omega}|j_{z}|$ for the Orszag–Tang vortex at $\mathrm{Re}=200\pi$ (magenta solid line), $\mathrm{Re}=2500$ (brown dashed line), and $\mathrm{Re}=5000$ (olive line with smaller dashes). The initial growth reflects large-scale magnetic field-line stretching, followed by a rapid intensification associated with current-sheet thinning and the onset of magnetic reconnection. Increasing Reynolds number leads to stronger current concentration and sustained intermittency, characteristic of turbulent MHD dynamics. Findings are obtained by the present RR scheme.
The energy evolution further clarifies the role of current-sheet dynamics in driving the non-linear MHD transition. The averaged kinetic and magnetic energies of the system are defined as
$E_{\mathrm{k}}=\frac{1}{2L^{2}}\int_{\Omega}|\mathbf{u}|^{2}\,\mathrm{d}\Omega,\quad E_{\mathrm{m}}=\frac{1}{2L^{2}}\int_{\Omega}|\mathbf{b}|^{2}\,\mathrm{d}\Omega,$ (37)
respectively. In all cases, the initial phase is characterised by a gradual amplification of magnetic energy through large-scale field-line stretching, while kinetic energy remains comparatively weak. As the system approaches the time of peak current density (see Fig. 4), a rapid drop in magnetic energy is observed simultaneously with a sharp rise in kinetic energy, indicating efficient conversion of magnetic energy into flow motion through reconnection events occurring within the thinned current sheets. This transfer becomes increasingly pronounced with Reynolds number: the turbulent regimes exhibit stronger magnetic energy depletion and enhanced kinetic energisation, consistent with the higher $j_{\max}$ values associated with more intense current-sheet collapse. Following the reconnection burst, both energies approach a slowly decaying, fluctuating state, reflecting sustained turbulent mixing and the continuous formation of secondary current sheets. In contrast, the low-Reynolds-number case rapidly relaxes toward a smoother configuration with significantly weaker energy exchange, in agreement with its lower peak current intensification.
Figure 4: Time evolution of kinetic (solid lines) and magnetic (dashed lines) energies for the Orszag–Tang vortex at $\mathrm{Re}=200\pi$ (orange), $\mathrm{Re}=2500$ (navy), and $\mathrm{Re}=5000$ (dark green). At early times, magnetic energy dominates as the initial field is stretched by the flow. As non-linear interactions intensify, magnetic energy is rapidly converted into kinetic energy through current-sheet formation and magnetic reconnection. Increasing Reynolds number leads to stronger energy transfer, delayed saturation, and enhanced turbulent mixing, with higher-Re cases exhibiting sustained kinetic activity and slower magnetic dissipation characteristic of fully developed MHD turbulence. Findings are obtained by the present RR scheme.
III.3 Computational cost
Figure 5: Computational cost: normalised runtime as a function of $N^{2}$ for the BGK (solid black line with squares), RMs (dashed red line with circles), CMs (green dotted line with triangles), and $R$ (blue dash-dotted line with inverted triangles) schemes in log-log scale. Moments-based LBMs consistently provide the most efficient performance.
We conclude the performance assessment of the lattice Boltzmann formulations for MHD with an evaluation of their computational cost. Fig. 5 summarises the normalised runtime of the four lattice Boltzmann models as a function of the total number of grid points, $N^{2}$. Analyses are carried out by varying $N\in[200:2000]$. While all approaches perform within a similar order of magnitude, clear and systematic differences emerge in the absolute runtime, reflecting the distinct algebraic complexity of the underlying collision operators.
Across the entire range of resolutions considered, the RR model displays the highest runtime. This behaviour is a direct consequence of the more elaborate collision step, which involves the projection of non-equilibrium populations onto Hermite polynomials followed by recursive reconstruction procedures. While these operations improve numerical stability, they introduce a measurable computational overhead that persists as the problem size increases.
In contrast, the RMs and CMs formulations consistently define the lower bound of the runtime envelope. The near-perfect overlap of their curves in Fig. 5 indicates that the additional velocity-dependent shift required by the CMs approach incurs a negligible cost relative to the dominant streaming operations. This result confirms that relaxation in moment space can be achieved without sacrificing computational efficiency.
Despite its apparent algorithmic simplicity, the BGK model does not achieve the lowest runtime. As shown in Fig. 5, both the RMs and CMs models outperform BGK across all tested resolutions. This suggests that the moment-space formulation enables a more efficient evaluation of the collision operator, due to reduced floating-point operations in the equilibrium reconstruction compared with the direct population-based BGK relaxation.
We can conclude that moment-based collision operators provide a favourable balance between computational efficiency and modelling sophistication, while the RR approach trades increased runtime for enhanced numerical robustness.
IV Conclusions
We have introduced and assessed a recursive regularised lattice Boltzmann method for two-dimensional incompressible magnetohydrodynamics. The proposed scheme retains the efficiency and locality of Dellar’s double-distribution framework for MHD, while enhancing the fluid solver through a Hermite-based recursive reconstruction of the non-equilibrium distribution. By combining a fourth-order Hermite equilibrium on D2Q9 with a regularised non-equilibrium built from physically consistent Hermite coefficients, the method selectively filters spurious lattice-supported kinetic contributions without requiring explicit velocity-gradient evaluations.
The Orszag–Tang vortex was used as a stringent benchmark to quantify accuracy, dissipation, and robustness against established collision models (BGK, raw-moment MRT, and central-moment formulations) under otherwise identical settings. In the well-resolved early evolution, all schemes closely match spectral reference data for peak current density and vorticity. As the flow develops sharp gradients and current sheets, the recursive regularisation introduces a modest additional dissipation on coarse grids, leading to slightly lower extrema compared with BGK. Importantly, this discrepancy diminishes under grid refinement: on the finest meshes, the RR results converge to the same level of agreement achieved by the competing collision operators, confirming consistency with the reference solution in the physically relevant regime.
We further examined the solenoidal constraint on the magnetic field. All tested lattice Boltzmann formulations display clear rapid decay of the maximum divergence error with increasing resolution, and the differences between collision models remain negligible. This indicates that the divergence behaviour is primarily controlled by the discrete magnetic-field representation (D2Q5) rather than by the regularisation of the fluid non-equilibrium moments. Qualitatively, the RR scheme reproduces the canonical Orszag–Tang dynamics, including the progressive distortion of large-scale magnetic structures, the formation of thin current sheets, and the emergence of intermittent filamentary features, in line with established MHD studies within the LBM community.
In turbulent regimes dominated by intermittent current sheets and reconnection-driven dynamics, recursive regularisation demonstrates robustness comparable to moment-based approaches while retaining the simplicity and efficiency of a single-relaxation-time framework. The accurate capture of peak current intensification, vorticity growth, and magnetic topology evolution confirms that the method preserves the essential physical mechanisms governing MHD turbulence.
From a computational standpoint, the recursive regularisation entails a moderate increase in per-time-step cost due to the additional Hermite projections and recursive reconstructions performed during the collision step. Nevertheless, the algorithm preserves the strictly local character of the lattice Boltzmann method and maintains linear scaling with system size. As a result, the added overhead remains controlled and does not compromise the favourable parallel efficiency or memory-access patterns of standard LBM formulations. In practical terms, the improved numerical robustness offered by the RR scheme can offset this additional cost by enabling stable simulations at coarser resolutions or under more demanding physical conditions.
In summary, our results demonstrate that recursive regularisation can be transferred to coupled MHD systems in a simple and systematic manner, delivering a robust stabilisation strategy while preserving the correct incompressible MHD limit. Beyond the benchmark considered here, the proposed formulation provides a practical route to regularised lattice Boltzmann solvers for broader multiphysics settings. Natural next steps include the extension to three-dimensional lattices foldes2023efficient; vasconez2026parametric and the development of consistent regularisation strategies for the magnetic populations to further reduce lattice artefacts at extreme transport parameters.
Supplemental material
The supplemental material is available at https://github.com/SIG-LBM-Multiphysics-Modelling/RecursiveRegularised_MHD. It includes all C++ programs used to reproduce the numerical simulations presented in this work. The implementations are written using the Kokkos library 9485033, ensuring performance portability across diverse hardware architectures.
Acknowledgements.
The author is grateful to Dr. Abinhav Muta for his help with the Kokkos library.